# Unboundedness for motivic invariants of birational automorphisms

### v1

@@ Abstract.
We introduce horizontal and vertical motivic invariants of birational maps between rational dominant maps and study their basic properties. As a first application, we show that the (usual) motivic invariants vanish for birational automorphisms of threefolds over algebraically closed fields of characteristic zero. On the other hand, we prove that the motivic invariants of the birational automorphism group of many types of varieties, including projective spaces of dimension at least four over a field of characteristic zero, do not form a bounded family, even after extending scalars to the algebraic closure of the field. For such varieties, we further show that their birational automorphism groups are not generated by maps preserving a conic bundle or a rational surface fibration structure, and their abelianizations do not stabilize.
@@ 1. Introduction
Here, ${\operatorname{ExDiv}}(\phi)$ is the set of exceptional divisors of $\phi$ and ${\operatorname{Burn}}_{*}(\mathbb{k})$ is the Burnside group [29], that is the free abelian group generated by the birational isomorphism classes of $\mathbb{k}$-varieties. This invariant has been generalized to birational maps betwen orbifolds in [30] and to volume preserving maps in [10, 37]. In this paper, we generalize the motivic invariant $c(\phi)$ to the relative setting and use it to prove new results about the groups of birational automorphisms.
@@ 1.1. Horizontal and vertical invariants
Let us explain the context of this theorem. The corresponding result for surfaces over an algebraically closed field is trivial, and over an arbitrary perfect field $\mathbb{k}$ the result is deduced from the Minimal Model Program and Sarkisov link factorization for surfaces in [35], where the study of such questions was initiated. In dimension $3$, the corresponding result is false [34, §3.3] over many nonclosed fields (such as $\mathbb{k}=\mathbb{Q}$), but it is known to be true when $\mathbb{k}$ is algebraically closed of characteristic zero and $X$ is rationally connected [34, Proposition 2.6]. In dimension at least $4$, the result is also false, even over $\mathbb{k}=\mathbb{C}$ [34, §3.4].
@@ 1.2. Applications
This result is new already in the case when $B=\mathbb{P}^{k}$, $k>0$, where it reproves and strengthens the known results that ${\operatorname{Cr}}_{n}(\mathbb{k})$ with $n\geq 4$ is not generated by linear automorphisms and de Jonquières maps [6, Theorem C] (which also holds for $m=3$), or by pseudo-regularizable elements [34, Theorem 1.2], [19, Theorem 1.2] (which is currently unknown for $m=3$). Our approach to proving such results is entirely different from Blanc–Lamy–Zimmermann [6] as we rely on motivic invariants while the proof of [6] is using Sarkisov link decomposition.
@@ 2.3. Rational Stein factorizations
    then the rational Stein factorization of $\pi$ is X ν^-1~π⇢~$B$ where $\nu\colon\widetilde{B}^{\nu}\to\widetilde{B}$ is the normalization of $\widetilde{B}$. In particular, if we additionally assume that $X$ is normal, then the rational Stein factorization of $\pi$ coincides with the usual one.
@@ 4.1. Unboundedness
Let $B$ be as in Lemma 4.4. There exists a sequence $\xi_{d}\colon J_{d}\to B$, $d\geq 1$ of nonisotrivial Weierstrass elliptic fibrations with the following properties:
@@
    The Mordell–Weil group of rational sections of $J_{d}\to B$ contains full $5$-torsion $(\mathbb{Z}/5)^{\oplus 2}$.
@@
The idea is to start with an appropriate Weierstrass fibration over $\mathbb{P}^{1}_{t}$ and pull it back to $B$ under the maps in (4.1). By [14, Theorem 5.1], there exists a elliptic curve $E/\mathbb{Q}(t)$ with nonconstant $j$-invariant and such that $(\mathbb{Z}/5)^{\oplus 2}\subset E(\mathbb{Q}(t))$. We make a scalar extension of $E/\mathbb{Q}(t)$ to $\mathbb{k}(t)$ and take a model $\xi^{\mathbb{P}^{1}}\colon J\to\mathbb{P}^{1}_{t}$. By construction the group of sections of $\xi^{\mathbb{P}^{1}}$ contains $(\mathbb{Z}/5)^{\oplus 2}$.
@@
Fix a Weierstrass model (see [40, Definition 1.1 and Theorem 2.1])
$\xi^{\mathbb{P}^{n}}\colon W(\mathscr{L},a,b)\to\mathbb{P}^{n},\text{ with }\mathscr{L}\in{\operatorname{Pic}}(\mathbb{P}^{n}),\;a\in\Gamma((\mathscr{L}^{\vee})^{\otimes 4}),\;\;b\in\Gamma((\mathscr{L}^{\vee})^{\otimes 6})$
birational to the pullback of $\xi^{\mathbb{P}^{1}}$ with respect to $p\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{1}$. Let $\xi_{d}^{\mathbb{P}^{n}}$ and $\xi_{d}\colon J_{d}\to B$ be the pullbacks of $\xi^{\mathbb{P}^{n}}$ under $F_{d}$ and $F_{d}\circ g$ respectively. By construction, these are Weierstrass elliptic fibrations that satisfy properties (2) and (3).
Since $\xi^{\mathbb{P}^{1}}$ is not isotrivial, so is $\xi^{\mathbb{P}^{n}}$, which implies $\mathscr{L}\not\simeq\mathscr{O}_{\mathbb{P}^{n}}$. As $a$ is a nonzero section of $(\mathscr{L}^{\vee})^{\otimes 4}$ (again because $\xi^{\mathbb{P}^{n}}$ is not isotrivial) the line bundle $\mathscr{L}^{\vee}$ is ample. We can assume that $d$ is large enough so that $\omega_{\mathbb{P}^{n}}\otimes F_{d}^{*}\mathscr{L}^{\vee}$ is ample. As $\xi_{d}^{\mathbb{P}^{n}}$ is the Weierstrass model
@@
if follows from the canonical bundle formula for Weierstrass fibrations [40, (1.2)(2)] that for such $d$, property (1) holds for $\xi_{d}^{\mathbb{P}^{n}}$, and hence also for $\xi_{d}$ because $g$ is finite.
@@ Appendix A Constructing elliptic torsors of prescribed prime index
Let $\mathbb{k}$ be a field of characteristic zero. Let $B$ be a geometrically integral $\mathbb{k}$-variety of dimension $n>0$. Let $p$ be a prime number and let $E$ be an elliptic curve over $\mathbb{k}(B)$ whose $j$-invariant is not in $\mathbb{k}$. Suppose that $(\mathbb{Z}/p)^{\oplus 2}\subset E(\mathbb{k}(B))$. Then there exist infinitely many $E$-torsors $\{C_{i}\}_{i\in\mathbb{N}}$ such that $C_{i,\overline{\mathbb{k}}(B)}$ are pairwise non isomorphic (as $\overline{\mathbb{k}}(B)$-varieties) and that each $C_{i,\overline{\mathbb{k}}(B)}$ has index $p$.
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\frac{E(\mathbb{k}(B))}{pE(\mathbb{k}(B))}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathbb{k}(B),E[p])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathbb{k}(B),E)[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\frac{E(\overline{\mathbb{k}}(B))}{pE(\overline{\mathbb{k}}(B))}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\overline{\mathbb{k}}(B),E[p])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\overline{\mathbb{k}}(B),E)[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
Since the $j$-invariant $j_{E}\in\mathbb{k}(B)$ of $E$ is not in $\mathbb{k}$ and $B$ is geometrically integral, so that $\overline{\mathbb{k}}\cap\mathbb{k}(B)=\mathbb{k}$, $j_{E}$ is not in $\overline{\mathbb{k}}$ neither. So both $\frac{E(\mathbb{k}(B))}{pE(\mathbb{k}(B))}$ and $\frac{E(\overline{\mathbb{k}}(B))}{pE(\overline{\mathbb{k}}(B))}$ are finite (see e.g. [13, Example 2.2]). As $E[p]\simeq(\mathbb{Z}/p)^{\oplus 2}$, we have
[…]
where $G_{\mathbb{F}}$ denotes the absolute Galois group of a field $\mathbb{F}$. By Lemma A.2, the image of the middle vertical map in the diagram is infinite. Hence the image of the composition
$H^{1}(\mathbb{k}(B),E[p])\to H^{1}(\overline{\mathbb{k}}(B),E[p])\to H^{1}(\overline{\mathbb{k}}(B),E)[p]$
is infinite, which gives rise to infinitely many $E$-torsors $C_{i}$ which are still non isomorphic as $E_{\overline{\mathbb{k}}(B)}$-torsors. There are only finitely many $E_{\overline{\mathbb{k}}(B)}$ torsor structures on a fixed curve of genus $1$ [43, Exercise 10.4], hence after removing repetitions we can assume that $C_{i,\overline{\mathbb{k}(B)}}$ are pairwise non isomorphic curves.
@@ Acknowledgments
We thank Jeremy Blanc, Michel Brion, Serge Cantat, Yuri Prokhorov, Pavel Sechin, Markus Szymik for discussions and their interest in our work. HYL is supported by the Yushan Fellow Program by the Ministry of Education (NTU-114V1006-5), the National Council for Science and Technology in Taiwan (114-2918-I-002 -020-, 114-2628-M-002 -012-), and the Asian Young Scientist Fellowship. Part of this work was done during his visit to Paris; he thanks Claire Voisin and IMJ-PRG for the hospitality, and the ERC grant HyperK (Grant agreement No. 854361) for its financial support. E.S. is supported by the UKRI Horizon Europe guarantee award ‘Motivic invariants and birational geometry of simple normal crossing degenerations’ EP/Z000955/1.

### v2

@@ Abstract.
We introduce horizontal and vertical motivic invariants of birational maps between rational dominant maps and study their basic properties. As a first application, we show that the (usual) motivic invariants vanish for birational automorphisms of threefolds over algebraically closed fields of characteristic zero. On the other hand, we prove that the motivic invariants of the birational automorphism groups of many types of varieties, including projective spaces of dimension at least four over a field of characteristic zero, do not form a bounded family, even after extending scalars to the algebraic closure of the field. For such varieties, we further show that their birational automorphism groups are not generated by maps preserving a conic bundle or a rational surface fibration structure, and their abelianizations do not stabilize.
@@ 1. Introduction
Here, ${\operatorname{ExDiv}}(\phi)$ is the set of exceptional divisors of $\phi$ and ${\operatorname{Burn}}_{*}(\mathbb{k})$ is the Burnside group [29], that is the free abelian group generated by the birational isomorphism classes of $\mathbb{k}$-varieties. This invariant has been generalized to birational maps between orbifolds in [30] and to volume preserving maps in [10, 37]. In this paper, we generalize the motivic invariant $c(\phi)$ to the relative setting and use it to prove new results about the groups of birational automorphisms.
@@ 1.1. Horizontal and vertical invariants
Let us explain the context of this theorem. The corresponding result for surfaces over an algebraically closed field is trivial, and over an arbitrary perfect field $\mathbb{k}$ the vanishing result for surfaces was deduced from the Minimal Model Program and Sarkisov link factorization in [35], where the study of such questions was initiated. In dimension $3$, the corresponding result is false [34, §3.3] over many nonclosed fields (such as $\mathbb{k}=\mathbb{Q}$), but it is known to be true when $\mathbb{k}$ is algebraically closed of characteristic zero and $X$ is rationally connected [34, Proposition 2.6]. In dimension at least $4$, the result is also false, even over $\mathbb{k}=\mathbb{C}$ [34, §3.4].
@@ 1.2. Applications
This result is new already in the case when $B=\mathbb{P}^{k}$, $k>0$, where it reproves and strengthens the known results that ${\operatorname{Cr}}_{n}(\mathbb{k})$ with $n\geq 4$ is not generated by linear automorphisms and de Jonquières maps [6, Theorem C] (which also holds for $n=3$), or by pseudo-regularizable elements [34, Theorem 1.2], [19, Theorem 1.2] (which is currently unknown for $n=3$). Our approach to proving such results is entirely different from Blanc–Lamy–Zimmermann [6] as we rely on motivic invariants while the proof of [6] is using Sarkisov link decomposition.
@@ 2.3. Rational Stein factorizations
    then the rational Stein factorization of $\pi$ is X ν-1~π~$B$ where $\nu\colon\widetilde{B}^{\nu}\to\widetilde{B}$ is the normalization of $\widetilde{B}$. In particular, if we additionally assume that $X$ is normal, then the rational Stein factorization of $\pi$ coincides with the usual one.
@@ 4.1. Unboundedness
Let $B$ be as in Lemma 4.4. There exists a sequence $\xi_{d}\colon J_{d}\to B$, $d\geq 1$ of nonisotrivial elliptic fibrations with the following properties:
@@
    The Mordell–Weil group of rational sections of $J_{d}\to B$ contains $5$-torsion subgroup $\mathbb{Z}/5$.
@@
The idea is to start with an appropriate elliptic fibration over $\mathbb{P}^{1}_{t}$ and pull it back to $B$ under the maps in (4.1). By [14, Theorem 5.1], there exists a elliptic curve $E/\mathbb{Q}(t)$ with nonconstant $j$-invariant and such that $\mathbb{Z}/5\subset E(\mathbb{Q}(t))$. We make a scalar extension of $E/\mathbb{Q}(t)$ to $\mathbb{k}(t)$ and take a model $\xi^{\mathbb{P}^{1}}\colon J\to\mathbb{P}^{1}_{t}$. By construction the group of sections of $\xi^{\mathbb{P}^{1}}$ contains $\mathbb{Z}/5$.
@@
Fix an elliptic fibration
$\xi^{\mathbb{P}^{n}}\colon\mathscr{E}\to\mathbb{P}^{n},$
birational to the pullback of $\xi^{\mathbb{P}^{1}}$ with respect to $p\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{1}$. Let $\xi_{d}^{\mathbb{P}^{n}}$ and $\xi_{d}\colon J_{d}\to B$ be the pullbacks of $\xi^{\mathbb{P}^{n}}$ under $F_{d}$ and $F_{d}\circ g$ respectively. By construction, these are elliptic fibrations that satisfy (3) and whose $j$-invariant map has Stein degree $\geq d$. The latter implies (2) once (1) is proved.
Now we prove (1). It suffices to prove the statement for $\xi_{d}\times_{\mathbb{k}}K$, where $K$ is an algebraically closed extension of $\mathbb{k}$ containing $\mathbb{C}$. Fix a Weierstrass model (see [40, Definition 1.1 and Theorem 2.1])
$\zeta^{\mathbb{P}^{n}}\colon W(\mathscr{L},a,b)\to\mathbb{P}^{n},\text{ with }\mathscr{L}\in{\operatorname{Pic}}(\mathbb{P}^{n}),\;a\in\Gamma((\mathscr{L}^{\vee})^{\otimes 4}),\;\;b\in\Gamma((\mathscr{L}^{\vee})^{\otimes 6})$
birational to $\xi^{\mathbb{P}^{n}}\times_{\mathbb{k}}K$. Let $\zeta_{d}^{\mathbb{P}^{n}}$ and $\zeta_{d}\colon J_{d}\to B$ be the pullbacks of $\zeta^{\mathbb{P}^{n}}$ under $F_{d}$ and $F_{d}\circ g$ respectively. Since $\xi^{\mathbb{P}^{1}}$ is not isotrivial, so is $\zeta^{\mathbb{P}^{n}}$, which implies $\mathscr{L}\not\simeq\mathscr{O}_{\mathbb{P}^{n}}$. As $a$ is a nonzero section of $(\mathscr{L}^{\vee})^{\otimes 4}$ (again because $\zeta^{\mathbb{P}^{n}}$ is not isotrivial) the line bundle $\mathscr{L}^{\vee}$ is ample. We can assume that $d$ is large enough so that $\omega_{\mathbb{P}^{n}}\otimes F_{d}^{*}\mathscr{L}^{\vee}$ is ample. As $\zeta_{d}^{\mathbb{P}^{n}}$ is the Weierstrass model
@@
if follows from the canonical bundle formula for Weierstrass fibrations [40, (1.2)(2)] that for such $d$, property (1) holds for $\zeta_{d}^{\mathbb{P}^{n}}$, and hence also for $\zeta_{d}$ because $g$ is finite. As $\xi_{d}\times_{\mathbb{k}}K$ is birational to $\zeta_{d}$, (1) is proved.
@@ Appendix A Constructing elliptic torsors of prescribed prime index
Let $\mathbb{k}$ be a field of characteristic zero. Let $B$ be a geometrically integral $\mathbb{k}$-variety of dimension $n>0$. Let $p$ be a prime number and let $E$ be an elliptic curve over $\mathbb{k}(B)$ whose $j$-invariant is not in $\mathbb{k}$. Suppose that $\mathbb{Z}/p\subset E(\mathbb{k}(B))$. Then there exist infinitely many $E$-torsors $\{C_{i}\}_{i\in\mathbb{N}}$ such that $C_{i,\overline{\mathbb{k}}(B)}$ are pairwise non isomorphic (as $\overline{\mathbb{k}}(B)$-varieties) and that each $C_{i,\overline{\mathbb{k}}(B)}$ has index $p$.
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\frac{E(\mathbb{k}(B))}{pE(\mathbb{k}(B))}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathbb{k}(B),E[p])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathbb{k}(B),E)[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\frac{E(\overline{\mathbb{k}}(B))}{pE(\overline{\mathbb{k}}(B))}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\overline{\mathbb{k}}(B),E[p])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${H^{1}(\overline{\mathbb{k}}(B),E)[p]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
Since the $j$-invariant $j_{E}\in\mathbb{k}(B)$ of $E$ is not in $\mathbb{k}$ and $B$ is geometrically integral, so that $\overline{\mathbb{k}}\cap\mathbb{k}(B)=\mathbb{k}$, $j_{E}$ is not in $\overline{\mathbb{k}}$ neither. So both $\frac{E(\mathbb{k}(B))}{pE(\mathbb{k}(B))}$ and $\frac{E(\overline{\mathbb{k}}(B))}{pE(\overline{\mathbb{k}}(B))}$ are finite (see e.g. [13, Example 2.2]).
Let $Q$ denote the quotient of $E[p]$ by $\mathbb{Z}/p$ as group schemes. Then we have a commutative diagram with exact rows:
${Q(\mathbb{k}(B))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\mathbb{k}(B),\mathbb{Z}/p)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta}$${H^{1}(\mathbb{k}(B),E[p])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Q(\overline{\mathbb{k}}(B))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\overline{\mathbb{k}}(B),\mathbb{Z}/p)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(\overline{\mathbb{k}}(B),E[p])}$
The vertical arrow in the middle is isomorphic to
${\operatorname{Hom}}(G_{\mathbb{k}(B)},\mathbb{Z}/p)\to{\operatorname{Hom}}(G_{\overline{\mathbb{k}}(B)},\mathbb{Z}/p),$
where $G_{\mathbb{F}}$ denotes the absolute Galois group of a field $\mathbb{F}$. By Lemma A.2, the image of this map is infinite. It follows that $\delta$ has infinite image as $Q(\overline{\mathbb{k}}(B))$ is finite. Since $\ker(\beta)\simeq\frac{E(\overline{\mathbb{k}}(B))}{pE(\overline{\mathbb{k}}(B))}$, which is finite, it follows that the image of the composition
${\operatorname{Hom}}(G_{\mathbb{k}(B)},\mathbb{Z}/p)\simeq H^{1}(\mathbb{k}(B),\mathbb{Z}/p)\overset{\delta}{\to}H^{1}(\overline{\mathbb{k}}(B),E[p])\xrightarrow{\beta}H^{1}(\overline{\mathbb{k}}(B),E)[p]$
is infinite, which gives rise to infinitely many $E$-torsors $C_{i}$ which are still non isomorphic as $E_{\overline{\mathbb{k}}(B)}$-torsors. There are only finitely many $E_{\overline{\mathbb{k}}(B)}$-torsor structures on a fixed curve of genus $1$ [43, Exercise 10.4], hence after removing repetitions we can assume that $C_{i,\overline{\mathbb{k}(B)}}$ are pairwise non isomorphic curves.
@@ Acknowledgments
We thank Jeremy Blanc, Michel Brion, Serge Cantat, Yuri Prokhorov, Pavel Sechin, Markus Szymik for discussions and their interest in our work. We thank an anonymous referee for pointing out a mistake in Proposition 4.5 in an earlier version of the paper. HYL is supported by the Yushan Fellow Program by the Ministry of Education (NTU-114V1006-5), the National Council for Science and Technology in Taiwan (114-2918-I-002 -020-, 114-2628-M-002 -012-), and the Asian Young Scientist Fellowship. Part of this work was done during his visit to Paris; he thanks Claire Voisin and IMJ-PRG for the hospitality, and the ERC grant HyperK (Grant agreement No. 854361) for its financial support. E.S. is supported by the UKRI Horizon Europe guarantee award ‘Motivic invariants and birational geometry of simple normal crossing degenerations’ EP/Z000955/1.